OpenStudy (anonymous):

what points is in the solution set of y < x^2 - 4x + 3? (2, -1)?

OpenStudy (lgbasallote):

simply substitute x = 2 and y = -1 -1 < (2)^2 - 4(-2) + 3 -1 < 4 + 8 + 3 -1 < 15 is this true? is -1 less than 15?

OpenStudy (anonymous):

-1 < 15 is right

OpenStudy (anonymous):

wait isnt it -16?

OpenStudy (lgbasallote):

4 +8 = 12 12 + 3 = 15

OpenStudy (anonymous):

-16<0 which is also true

OpenStudy (lgbasallote):

anyway since it's true that -1 < 15 then you are right

OpenStudy (anonymous):

ok thanks.

OpenStudy (lgbasallote):

wait you're right...i made a mistake in my calculation -1< (2)^2 - 4(2) + 3 -1 < 4 - 8 + 3 -1 < -1 this is not true so it's not a solution

OpenStudy (anonymous):

okay, so i got it wrong the first time yeah?

OpenStudy (anonymous):

if i did, i shall try again..

OpenStudy (lgbasallote):

yes. you need to try another ordered pair

OpenStudy (anonymous):

okay, can you check my answer again after i show my work? thanks.

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

i tried (-3, 0) -3 < (0)2 - 4(0) + 3 =-6

OpenStudy (anonymous):

i think this is true, do you think so?>

OpenStudy (lgbasallote):

-3 < (0)^2 - 4(0) + 3 -3 < 3 yes it is true

OpenStudy (lgbasallote):

however...

OpenStudy (lgbasallote):

you substituted it wrong

OpenStudy (lgbasallote):

should be 0 < (3)^2 - 4(3) + 3 0< 9 - 12 + 3 0 < 0 this is not true

OpenStudy (anonymous):

hmm, so its wrong again.. aish:( ok i'll try it again

OpenStudy (lgbasallote):

go for it

OpenStudy (anonymous):

i tried (1, 0) so 0 < (1)2 - 4(1) + 3 0 <1x2-4x1+3 =-1?

OpenStudy (lgbasallote):

it's not 1 x 2 it's 1^2 that means 1 x 1

OpenStudy (anonymous):

ooh oops sorry i'll try i tagin.

OpenStudy (anonymous):

again*

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

0 < 1^2 - 4(0) + 3 =-4?

OpenStudy (lgbasallote):

why did it suddenly become 4(0)?

OpenStudy (anonymous):

oh God. i get confused. so sorry! wwait.

OpenStudy (anonymous):

0 < 1^2 - 4(1) + 3

OpenStudy (lgbasallote):

right. what does that become then?

OpenStudy (anonymous):

0<0 false

OpenStudy (lgbasallote):

correct

OpenStudy (lgbasallote):

so find a different ordered pair

OpenStudy (anonymous):

i think (-3, 0) is right.. 0 < (-3)2 - 4(-3) + 3 =-9, true.

OpenStudy (lgbasallote):

you're 1/3 right and 3/4 wrong hehe you're 2/3 wrong because 1) 0 is NOT less than -9 2) (-3)^2 - 4(-3) + 3 is NOT equal to -9 you're 1/3 right because 0 < (-3)^2 - 4(-3) + 3 0 < 9 + 12 + 3 0 < 24 and this is true

OpenStudy (lgbasallote):

you're 1/3 right and 2/3 wrong* my math was wrong in my first statement =))

OpenStudy (anonymous):

i dont get it.. :(

OpenStudy (lgbasallote):

you did a wrong calculation

OpenStudy (lgbasallote):

0 < (-3)^2 - 4(-3) + 3 0 < 9 + 12 + 3 0 < 21 + 3 0 < 24 thus his is true

OpenStudy (lgbasallote):

thus (-3,0) is correct

OpenStudy (anonymous):

:/ oh.. okay. i got it not. thank you!!

OpenStudy (anonymous):

got it now*

OpenStudy (lgbasallote):

welcome^_^